OpenStudy (anonymous):

How do you simplify 153/34?

OpenStudy (anonymous):

4.5, if you can use a calculator, just plug and chug those numbers.

OpenStudy (anonymous):

I don't get what you're saying, but thanks.

OpenStudy (anonymous):

If you would like for me to explain something further please let me know. I don't mind helping.

OpenStudy (anonymous):

9/2 I got it now.. thanks

OpenStudy (anonymous):

153/34 = 9/2 = 4.5